OpenStudy (anonymous):

Currently, Sam's coach is twice as old as Sam. When Sam started soccer 10 years ago, his coach was four times his age. How old is Sam now? 10 25 20 15

OpenStudy (anonymous):

Coach is 30 right now. Sam is 15 since half of 30 is 15. Now if you take away 10 from 15 you get that sam was 5 years old. Now to make sure it is correct you multiply 5 times 4 and you get 20 years. Now 30-20 is ten years. So Sam is now 15.

OpenStudy (anonymous):

let the coach be c and sam s c = 2s 4(s-10) = c-10 solve the system

OpenStudy (anonymous):

4s-40 = c-10, BUT c=2s 4s-40 = 2s-10 2s = 30 s = 15 c = 2(15) = 30 ^that's the math sans guess and check